### v2

@@ 7 Conclusion
In this paper, we analyzed the key factors in existing LLM evaluation frameworks that may lead to unreliable assessment results, highlighting the critical answer extraction module in LLM responses as a weak point. To address this issue, we introduce a key answer extraction model, xFinder, specifically for extracting key answers in LLM responses. Extensive experimental results demonstrate that xFinder, compared to existing RegEx-based key answer extraction methods, can extract key answers from LLM outputs more accurately and efficiently, thereby improving evaluation reliability. This study marks the first step in constructing a reliable evaluation framework. In the future, we will continue addressing other key issues in the reliable evaluation of LLMs, providing a foundation for accurately evaluating LLM performance.
@@ B.1.1 Training Set
The training set contains 26,900 samples. Detailed information about the models, key answer types, and prompt configurations used in the training set is presented in Tables  5,  6,  7, and  8, respectively.
@@
Table 7: Number of samples for each key answer type in the training set.
@@ B.1.2 Test Set
The test set comprises 4,961 samples. Detailed information about the models, key answer types, and prompt configurations used in the test set is presented in Tables  9,  10,  11, and  12, respectively.
@@
Table 9: Number of samples from each LLM in the test set.
@@
Table 10: Number of samples from each dataset in the test set.
@@
Table 11: Number of samples for each key answer type in the test set.
@@ B.1.3 Generalization Set
The generalization set comprises 4,482 samples. Detailed information about the models, key answer types, and prompt configurations used in the generalization set is presented in Tables  13,  14,  15, and  16, respectively.
@@
Table 13: Number of samples from each LLM in the generalization set.
@@ B.2 Human Annotation Details
In our manual annotation project, all annotators are volunteers with a background in natural language processing, and each annotator holds at least a master’s degree. The project team consists of five annotators, with a gender ratio of three men to two women. Regarding their remuneration, they first receive the standard employee salary. We also pay them an additional 2 RMB per annotated data item, with the average time for annotating each task being about 20 seconds. Additionally, our engineering team participates in data review, primarily responsible for supervising and evaluating annotation quality. The entire annotation process lasted for 15 days.
@@ Appendix C Fine-Tuning Hyperparameters
Table 17: xFinder fine-tuning hyperparameter setting.
@@ Appendix D Detailed Experimental Results
D.1 Extraction Accuracy Results
@@
alphabet option short text categorical label math Overall
@@
alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
@@
D.2 Evaluation in Real-World Scenarios Results
D.2.1 Alphabet & Short Text Option Tasks
@@
Table 20: Evaluation results of LLMs on the ARC-c (alphabet) task, presenting accuracy scores using different key answer extraction methods.
@@
D.2.2 Categorical Label Option Tasks
@@
D.2.3 Math Option Tasks
@@ F.1 Prompts for LLM Responses Generation & Real-World Scenarios Evaluation
Figure 19: 0-shot prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 20: 0-shot-restrict prompt for LLM response generation & Real-World scenarios evaluation (using alpha option type question as an example).
@@
Figure 21: 0-shot-cot prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 22: 0-shot-cot-restrict prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 23: 5-shot prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 24: 5-shot-restrict prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 25: 5-shot-cot prompt for LLM response generation (using alpha option type question as an example).
@@
Figure 26: 5-shot-cot-restrict prompt for LLM response generation (using alpha option type question as an example).
@@ F.2 Prompts for Auto Labeling Using GPT-4
You are a help assistant tasked with extracting the precise key answer from given output sentences. You must only provide the extracted key answer without including any additional text. I will provide you with a question, output sentences along with an answer range. The output sentences are the response of the question provided. The answer range could either describe the type of answer expected or list all possible valid answers. Using the information provided, you must accurately and precisely determine and extract the intended key answer from the output sentences. Please don’t have your subjective thoughts about the question. If the output sentences present multiple different answers, carefully determine if the later provided answer is a correction or modification of a previous one. If so, extract this corrected or modified answer as the final response. Conversely, if the output sentences fluctuate between multiple answers without a clear final answer, you should output [No valid answer]. For a question of type "classify_text", if the output sentences do not clearly provide any answer within the specified answer range, then the output should be [No valid answer]. Additionally, if the answer range is a list and the answer given in the output sentences is not included in the answer range, also output [No valid answer]. <?xml version="1.0" ?> <information> <question type="str">{question}</question> <output_sentences type="str">{LLM response}</output_sentences> <key_answer_type type="str">{key answer type}</key_answer_type> <answer_range type="dict"> <type type="str">a list of all possible valid answers</type> <choices type="list">{answer range}</choices> </answer_range> </information> Key extracted answer:
@@
F.3 Prompts for Extracting Key Answers Using xFinder

### v3

@@ 6 Conclusion
In this paper, we analyzed key issues in existing LLM evaluation frameworks that could result in unreliable outcomes, with a particular focus on the extraction and matching of LLM responses. In response to these challenges, we introduced xFinder, a novel evaluator specifically designed for accurate key answer extraction and matching in LLM responses. Extensive experiments showed that xFinder extracted key answers from LLM outputs more accurately and efficiently than existing RegEx-based methods, significantly improving evaluation reliability. This study represents the first step in constructing a reliable evaluation framework. In the future, we will have continued addressing other key issues in the reliable evaluation of LLMs, providing a foundation for accurately evaluating LLM performance.
@@ B.1.1 Training Set
The training set contains 26,900 samples. Detailed information about the models, key answer types, and prompt configurations used in the training set is presented in Tables  9,  10,  11, and  12, respectively.
@@
Table 11: Number of samples for each key answer type in the training set.
@@ B.1.2 Test Set
The test set comprises 4,961 samples. Detailed information about the models, key answer types, and prompt configurations used in the test set is presented in Tables  13,  14,  15, and  16, respectively.
Table 13: Number of samples from each LLM in the test set.
@@
Table 14: Number of samples from each dataset in the test set.
@@
Table 15: Number of samples for each key answer type in the test set.
@@ B.1.3 Generalization Set
The generalization set comprises 4,482 samples. Detailed information about the models, key answer types, and prompt configurations used in the generalization set is presented in Tables  17,  18,  19, and  20, respectively.
Table 17: Number of samples from each LLM in the generalization set.
@@ B.2 Human Annotation Details
In our manual annotation project, all annotators are volunteers with a background in natural language processing, and each annotator holds at least a master’s degree. The project team consists of five annotators, with a gender ratio of three men to two women. Regarding their remuneration, they first receive the standard employee salary. We also pay them an additional 2 RMB per annotated data item, with the average time for annotating each task being about 20 seconds. Additionally, our engineering team participates in data review, primarily responsible for supervising and evaluating annotation qualityEvery data item is annotated by two different annotators. We will recheck the annotation and make the final decision if two annotators have different results towards the same item The entire annotation process lasted for 15 days. Below are the annotation guidelines.
B.2.1 Human Annotation Guidelines
The annotation interface is shown in Figure 5. Each sample consists of the following elements: Question (the question), LLM Output (the response provided by a LLM for this question), Key answer type (the type of answer for this question, which can be one of: alphanumeric_numeric_options / short_text_content / classify_text / numerical), and Answer range (the set of candidate answers for this question).
Figure 5: Label Studio Interface.
Your Task. You are required to fill in the following two blanks.
  • The first blank is for the answer extracted from the LLM Output. For non-mathematical questions, the extracted answer MUST exactly match one of the candidate answers in the Answer range. For mathematical questions, directly extract the final answer from the LLM Output, preserving the original format (including any LaTeX formatting). When no valid answer can be extracted from the LLM Output (e.g., if the answer is irrelevant or inconsistent), enter [No valid answer] (without quotation marks).
  • The second blank is generally left empty, but can be used for notes in special cases. If you encounter a particularly unusual LLM Output, you may leave a comment here. In other cases, leave this field blank.
Annotation Details.
For alphanumeric option questions (alphanumeric_numeric_options):
  • The LLM Output may provide both the option letter and the corresponding option content, but you should only extract the option letter, i.e., “A/B/C/D” (note that the letter should be uppercase, with no additional spaces, and without quotation marks).
  • If the LLM Output provides both the option letter and content, but they do not match, enter [No valid answer] (without quotation marks).
For short text questions (short_text_content):
  • If the LLM Output correctly provides one of the candidate short texts from the Answer range as the answer, label it with that short text.
  • Note that some candidate short texts in the Answer range may include punctuation (e.g., “.”), which must be copied and pasted exactly as the final label, without quotation marks.
For text classification questions (classify_text):
  • Similarly, only if the LLM Output clearly provides a candidate answer from the Answer range should you label it with that answer; otherwise, label it as [No valid answer] (without quotation marks).
  • If the answer in the LLM Output has some kind of relationship with a candidate answer in the Answer range (e.g., inclusion, intersection), but does not explicitly indicate that candidate answer, label it as [No valid answer]. For example, if the LLM Output is “The answer is River.” and the Answer range is [“Nature”, “Technology”], it should be labeled as [No valid answer].
For mathematical questions (numerical):
  • If the LLM Output does not explicitly provide an answer to the corresponding question, label it as [No valid answer].
  • If the LLM Output initially gives an answer ans1, but through detailed calculations, it rejects ans1 and provides a final answer ans2, extract ans2 as the final answer.
  • The label should not include commas or non-LaTeX symbols (such as “\$”). For instance, if the LLM Output is “The answer is \$10,000.”, the label should be “10000” (without quotation marks). But, the “\$” symbol with the LaTex equation should be retained.
B.3 Examples from the KAF Dataset
"key_answer_type": "alphabet option",
"question": "A man is seen playing guitar on a stage with others playing instruments behind him. The man grabs a guitar from the audience and begins playing both one after the other ...",
"llm_output": "Option A is the correct choice as it describes ...",
"standard_answer_range": "[[’A’, strums the guitar in the end, continues playing the guitar with the crowd following him as well as lining up next to him.’], [’B’, continues playing the instruments and ends by waving to the crowd and walking off stage.’], [’C’, then turns to the audience and gives a stuffed toy to the audience and continues playing.’], [’D’, finally stops playing and moves his hands for the crowd to see.’]]",
"gold_label": "A",
"xFinder_output": "A",
"key_answer_type": "short text",
"question": "If you really wanted a grape, where would you go to get it? Answer Choices: winery / fruit stand / field / kitchen / food",
"llm_output": "The answer is winery / fruit stand / field / kitchen / food ...",
"standard_answer_range": "[\"winery\", \"fruit stand\", \"field\", \"kitchen\", \"food\"]",
"gold_label": "[No valid answer]",
"xFinder_output": "[No valid answer]",
"key_answer_type": "categorical label",
"question": "How tall is the Sears Building?",
"llm_output": "The Sears Building is a specific structure, so the answer would be a Location ...",
"standard_answer_range": "[’Abbreviation’, Entity’, Description’, Person’, Location’, Number’]",
"gold_label": "Location",
"xFinder_output": "Location",
"key_answer_type": "math",
"question": "Mike made 69 dollars mowing lawns over the summer. If he spent 24 dollars buying new mower blades, how many 5 dollar games could he buy with the money he had left?",
[…]
"standard_answer_range": "a(n) number / set / vector / matrix / interval / expression / function / equation / inequality",
"gold_label": "9",
"xFinder_output": "9",
@@ Appendix C Fine-Tuning Hyperparameters
Table 21: xFinder fine-tuning hyperparameter setting.
@@ Appendix D Detailed Experimental Results
D.1 Example of Extraction vs. Judgment Accuracy
For the concepts of “Judgment Accuracy” and “Extraction Accuracy” mentioned in the main text, although we have provided explanations in the corresponding table notes, these concepts might still be somewhat difficult to understand. Therefore, we present a real evaluation example by OpenCompass to help clarify these concepts.
"question": "Which pair dont reproduce the same way? Answer Choices: (A) ... (B) ... (C) cat and catfish (D) caterpillar and butterfly",
"correct_answer": "C",
"tested_llm_output": "(A) is not the answer... So the correct answer is (D) Caterpillar and butterfly ...",
"correct_key_answer": "D",
"is_tested_llm_output_right": "Wrong",
Referring to the test data above, OpenCompass’s extraction is “A”, which is inconsistent with the “correct answer”, so OpenCompass’s judgment of the LLM’s output is “Wrong”. Comparing OpenCompass’s extraction (“A”) with the “correct key answer” (“D”), we can conclude that OpenCompass’s extraction is incorrect, which relates to extraction accuracy. Similarly, OpenCompass’s judgment (“Wrong”) aligns with the “Is the tested LLM output right?” result (“Wrong”), meaning that OpenCompass’s judgment is correct, which relates to judgment accuracy.
A 100% reliable evaluation logic first extracts the key answer and then compares it with the correct answer to make a judgment. The comparison between the key answer and the correct answer is a simple regular expression match, ensuring no errors. Under this assumption, correct extraction is a necessary and sufficient condition for correct judgment. However, current evaluation methods are not 100% reliable, which means that even if an evaluation method fails to correctly extract the key answer, it may still “guess” the final evaluation result correctly. This directly leads to the judgment accuracy of the same evaluation method being higher than its extraction accuracy. The fundamental reason for the discrepancy between the two metrics is extraction errors.
D.2 Extraction Accuracy Results
@@
xFinder alphabet option short text categorical label math Overall
@@
xFinder alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
@@
D.3 Evaluation in Real-World Scenarios Results
D.3.1 Alphabet & Short Text Option Tasks
Table 24: Evaluating results of LLMs on the ARC-c (alphabet) task, presenting accuracy scores using different key answer extraction methods.
@@ D.3.1 Alphabet & Short Text Option Tasks
D.3.2 Categorical Label Option Tasks
@@
D.3.3 Math Option Tasks
@@ F.1 Prompts for LLM Responses Generation & Real-World Scenarios Evaluation
Figure 20: 0-shot prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 21: 0-shot-restrict prompt for LLM response generation & Real-World scenarios evaluation (using alphabet option type question as an example).
@@
Figure 22: 0-shot-cot prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 23: 0-shot-cot-restrict prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 24: 5-shot prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 25: 5-shot-restrict prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 26: 5-shot-cot prompt for LLM response generation (using alphabet option type question as an example).
@@
Figure 27: 5-shot-cot-restrict prompt for LLM response generation (using alphabet option type question as an example).
@@ F.2 Prompts for Auto Labeling Using GPT-4
You are a help assistant tasked with extracting the precise key answer from given output sentences. You must only provide the extracted key answer without including any additional text. I will provide you with a question, output sentences along with an answer range. The output sentences are the responses to the question provided. The answer range could either describe the type of answer expected or list all possible valid answers. Using the information provided, you must accurately and precisely determine and extract the intended key answer from the output sentences. Please don’t have your subjective thoughts about the question. If the output sentences present multiple different answers, carefully determine if the later provided answer is a correction or modification of a previous one. If so, extract this corrected or modified answer as the final response. Conversely, if the output sentences fluctuate between multiple answers without a clear final answer, you should output [No valid answer]. For a question of type "classify_text", if the output sentences do not clearly provide any answer within the specified answer range, then the output should be [No valid answer]. Additionally, if the answer range is a list and the answer given in the output sentences is not included in the answer range, also output [No valid answer]. <?xml version="1.0" ?> <information> <question type="str">{question}</question> <output_sentences type="str">{LLM response}</output_sentences> <key_answer_type type="str">{key answer type}</key_answer_type> <answer_range type="dict"> <type type="str">a list of all possible valid answers</type> <choices type="list">{answer range}</choices> </answer_range> </information> Key extracted answer:
@@
F.3 Prompts for Judge Models Using GPT-4
As shown in Figure 30, this is the prompt used by GPT-4 as Judge in our experiments. This prompt is also applied to other judge models used for comparison in our study. Figure 31 illustrates the CoT prompt used by GPT-4 as Judge (CoT).
You are a diligent and precise assistant tasked with evaluating the correctness of responses. We request your feedback on whether the model’s response correctly answers the user question above. Based on the reference answer, please determine if the model’s answer is correct. Please first output "correct" or "incorrect" on a single line. In the subsequent line, provide a brief explanation of your judgment, ensuring it is objective and clear. Question: {question} Reference Answer: {reference} Model’s Answer: {answer}
Figure 30: The prompt for GPT-4 as Judge.
You are a diligent and precise assistant tasked with evaluating the correctness of responses. Think step by step as you make your evaluation. We request your feedback on whether the model’s response correctly answers the user question above. Follow these steps to make your evaluation: 1. Think step by step: Read the user question carefully. 2. Think step by step: Review the reference answer and understand the key points it covers. 3. Think step by step: Compare the model’s answer with the reference answer. 4. Think step by step: Determine if the model’s answer addresses the key points in the reference answer and correctly answers the question. First, provide your reasoning in detail. Then, clearly state your judgement as either "Correct" or "Incorrect." Please present your response in the following JSON format: {{ "reasoning": "Your step-by-step reasoning here.", "judgement": "Correct or Incorrect" }} Question: {question} Reference Answer: {reference} Model’s Answer: {answer}
Figure 31: The prompt for GPT-4 as Judge (CoT).
F.4 Prompts for Extracting Key Answers Using xFinder